### v2

# Constraining the nonstandard propagating gravitational waves in the cosmological background with GWTC-3

Abstract
The detection of gravitational waves (GWs) has opened a new window to test the fundamental nature of gravity. We present constraints on the nonstandard propagation of GWs using the spectral siren method applied to binary black hole (BBH) mergers from the third Gravitational-Wave Transient Catalog (GWTC-3). The spectral siren method exploits the redshift distribution of BBHs to probe the cosmic expansion history and break degeneracies between cosmology and modified gravity effects. We focus on the friction term $\nu$ in the nonstandard GW propagation equation, which characterizes the running of the Planck mass. Assuming the standard $\Lambda$CDM cosmology, we find $\nu=0.5^{+3.5}_{-2.6}$ (median and $90\%$ credible interval), improving upon previous constraints from the bright siren event GW170817 by an order of magnitude. This improvement is due to the higher redshifts of BBHs in GWTC-3, reaching up to $z\sim 1$. Our result suggests that the propagation of GWs is consistent with the predictions of general relativity, placing limits on modified gravity theories that predict a time-varying Planck mass. As the sensitivity of GW detectors improves, the spectral siren method will provide a powerful tool for testing gravity on cosmological scales and probing the physics of the early Universe.
Introduction. The quest to understand the fundamental nature of gravity and its role in the Universe has led to the exploration of modified gravity theories. Two primary motivations drive the pursuit of these alternative theories: the need to explain cosmic acceleration and the desire to find a substitute for dark matter. The observed accelerated expansion of the Universe [1, 2] and the presence of dark matter [3, 4] have challenged the standard model of cosmology based on general relativity (GR) and the $\Lambda$CDM paradigm. Modified gravity theories aim to address these challenges by introducing additional degrees of freedom or modifying the gravitational action [5, 6, 7, 8, 9]. These modifications can manifest in various ways, including altering the propagation of gravitational waves (GWs) through the cosmological background [10, 11, 12].
Modified gravity theories typically introduce changes in the weak-field regime at large scales, while reducing to GR in the strong-field regime through screening mechanisms such as Chameleon [13, 14], Vainshtein [15, 16], or Symmetron [17, 18]. These screening mechanisms ensure that the theories remain consistent with local tests of gravity [19, 20]. Cosmological tests of modified gravity often focus on the propagation of GWs rather than their generation [10, 11, 12]. Even if the modification to gravity is a tiny effect, the propagation of GWs over long distances can accumulate the effect, making it detectable by current and future GW detectors [21, 22].
The detection of GWs by the Advanced LIGO and Advanced Virgo detectors [23, 24, 25, 26, 27] has opened up a new observational window to test modified gravity theories, complementing the existing probes based on electromagnetic waves. One powerful method to constrain these theories is the bright siren approach, which relies on the detection of GWs with associated electromagnetic counterparts [28, 29, 30, 31]. A notable example of a bright siren is the binary neutron star merger event GW170817 with its accompanying kilonova emission [24, 32]. However, the applicability of the bright siren method is limited by the rarity of such events and the need for precise sky localization.
An alternative approach is the “spectral siren” method, which exploits features in the mass distribution of compact binaries, such as gaps or peaks, to constrain their redshift distribution [33]. By tracking the redshifting of these features, one can probe the expansion history independently of electromagnetic observations. While spectral sirens may not measure the Hubble constant as precisely as bright sirens [34], they have the potential to significantly outperform them in constraining departures from GR in GW propagation, as we will demonstrate in this letter.
In this letter, we focus on how nonstandard propagation of GWs can be constrained using the spectral siren method. Modified gravity theories often predict a departure from the standard GR wave equation, introducing additional terms that can be interpreted as a friction effect [10, 11, 35]. This modification accumulates over the vast distances traveled by cosmological GWs, making their propagation a sensitive probe of even small deviations from GR [11, 12, 36]. We aim to constrain these nonstandard propagation effects using the third GW Transient Catalog (GWTC-3) [37], leveraging the spectral siren method to break degeneracies between cosmology and modified gravity. Remarkably, we find that spectral sirens can constrain the friction term in modified GW propagation an order of magnitude better than bright sirens, highlighting their complementary role in probing gravity.
Theoretical Framework for Nonstandard GW Propagation. The propagation of GWs in the cosmological background can be modified by alternative theories of gravity, leading to observable deviations from the predictions of GR. A generalized framework has been developed to describe and test such nonstandard GW propagation, based on the effective field theory approach [11, 38, 39].
In this framework, the propagation equation for GWs is given by
$h_{ij}^{\prime\prime}+(2+\nu)\mathcal{H}h_{ij}^{\prime}+(c_{\mathrm{T}}^{2}k^{2}+a^{2}\mu^{2})h_{ij}=a^{2}\Gamma\gamma_{ij},$ (1)
where $h_{ij}$ is the metric perturbation, $\mathcal{H}\equiv a^{\prime}/a$ is the conformal Hubble parameter, and the prime denotes the derivative with respect to conformal time. The equation contains four time-dependent parameters $(\nu,c_{\mathrm{T}},\mu,\Gamma)$ that characterize possible deviations from GR:
  • $\nu$ is related to the time variation of the gravitational coupling strength, arising from the running of the Planck mass [40].
  • $c_{\mathrm{T}}$ is the propagation speed of GWs, which can differ from the speed of light.
  • $\mu$ represents the effective mass of the graviton.
  • $\Gamma\gamma_{ij}$ is a source term due to anisotropic stress, which can modify the GW amplitude.
In the limit where $\nu=0$, $c_{\mathrm{T}}=1$, $\mu=0$, and $\Gamma=0$, the propagation equation reduces to the standard GR case:
$h_{ij}^{\prime\prime}+2\mathcal{H}h_{ij}^{\prime}+k^{2}h_{ij}=0.$ (2)
This shows that GR is a special case within the generalized framework, corresponding to the absence of any modification parameters.
In this letter, we focus on the effects of the $\nu$ parameter, which is the least constrained by current observations. The GW170817 event has placed a stringent bound on the speed of GWs, $|c_{\mathrm{T}}-1|\lesssim 10^{-15}$ [41], allowing us to set $c_{\mathrm{T}}=1$. The graviton mass $\mu$ is also tightly constrained by various observations, such as solar system tests and binary pulsar observations [42], justifying the assumption $\mu=0$. The source term $\Gamma$ is neglected for simplicity, as it is expected to be subdominant in most scenarios. Under these assumptions, the propagation equation simplifies to
$h_{ij}^{\prime\prime}+(2+\nu)\mathcal{H}h_{ij}^{\prime}+k^{2}h_{ij}=0.$ (3)
This equation describes the nonstandard propagation of GWs due to the running of the Planck mass, characterized by the $\nu$ parameter [11].
Assuming $\nu$ varies slowly over cosmological timescales, while the GW wavelength is much smaller than the horizon scale, the WKB solution to the propagation equation can be expressed as [11]
$h=C_{\mathrm{MG}}h_{\mathrm{GR}},$ (4)
where $h_{\mathrm{GR}}$ is the GR waveform and the correction factor $C_{\mathrm{MG}}$ is given by
$C_{\mathrm{MG}}=e^{-\mathcal{D}},$ (5)
with $\mathcal{D}$ quantifying the modification to the GW amplitude:
$\mathcal{D}=\frac{1}{2}\int_{0}^{z}\frac{\nu}{1+z^{\prime}}\mathrm{d}z^{\prime}.$ (6)
The luminosity distance in the presence of nonstandard GW propagation can be derived by considering the energy flux of GWs. In GR, the luminosity distance $d_{L}^{\mathrm{GR}}$ is related to the redshift $z$ by [43]
$d_{L}^{\mathrm{GR}}(z)=(1+z)\int_{0}^{z}\frac{c\,\mathrm{d}z^{\prime}}{H(z^{\prime})},$ (7)
where $c$ is the speed of light and $H(z)$ is the Hubble parameter as a function of redshift.
In the nonstandard propagation scenario, the GW amplitude is modified by the factor $C_{\mathrm{MG}}$, which leads to a modified luminosity distance $d_{L}^{\mathrm{MG}}$ [22, 44]:
$d_{L}^{\mathrm{MG}}(z)=\frac{d_{L}^{\mathrm{GR}}(z)}{C_{\mathrm{MG}}(z)}=d_{L}^{\mathrm{GR}}(z)e^{\mathcal{D}(z)}.$ (8)
For the purposes of this paper, we assume that the friction term $\nu$ remains constant within the observable redshift range (i.e., for $z\lesssim 2$), following the approach of [45]. Combining the expressions for $d_{L}^{\mathrm{GR}}(z)$ and $\mathcal{D}(z)$, we obtain the relation between the modified luminosity distance and redshift [11, 46]:
$d_{L}^{\mathrm{MG}}(z)=(1+z)^{\nu/2}d_{\mathrm{GR}}$ (9)
where
$d_{\mathrm{GR}}=\frac{(1+z)}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{\sqrt{\Omega_{\mathrm{m}}(1+z^{\prime})^{3}+(1-\Omega_{\mathrm{m}})}}.$ (10)
Here, $H_{0}$ denotes the Hubble constant and $\Omega_{\mathrm{m}}$ is the matter density parameter at present. Besides, $(1-\Omega_{\mathrm{m}})$ represents the fraction of the Universe’s energy density in the form of dark energy, assuming a flat Universe and neglecting the contribution of radiation.
Eq. (8) shows that the nonstandard propagation of GWs, characterized by the $\nu$ parameter, leads to a modification of the luminosity distance-redshift relation compared to the GR case. By measuring the luminosity distances of GW events at different redshifts, it is possible to constrain the $\nu$ parameter and test the validity of GR on cosmological scales [22, 44, 47, 48].
This simplified framework allows us to isolate and constrain the effects of the $\nu$ parameter on the propagation of GWs using the GWTC-3 catalog from the LIGO-Virgo-KAGRA Collaboration [27]. By comparing the observed GW signals with the predictions of this nonstandard propagation model, we can place bounds on the possible time variation of the gravitational coupling strength, providing a test of GR on cosmological scales.
Methodology. We utilize a set of 42 binary black holes (BBHs) from the GWTC-3 catalog [27], selected based on a network-matched filter signal-to-noise ratio threshold of 11 and an inverse false alarm rate higher than 4 years, as outlined in [34, 49]. Our analyses incorporate combined posterior samples derived from the IMRPhenom [50, 51] and SEOBNR [52, 53] waveform families.
GW experiments measure the luminosity distance $D_{\mathrm{L}}$ and redshifted masses $m_{1}^{\mathrm{det}},m_{2}^{\mathrm{det}}$ for each BBH event, rather than the source redshift $z$ and intrinsic masses $m_{1}$, $m_{2}$. These quantities are connected via the relation:
$m_{i}=\frac{m_{i}^{\text{det}}}{1+z(D_{\mathrm{L}}^{\mathrm{MG}};H_{0},\Omega_{\mathrm{m}})}.$ (11)
By characterizing the source mass distribution, this relation enables probing the cosmic expansion history without relying on redshift information from electromagnetic counterparts [54, 55].
Given the observed data $\textbf{d}=\{d_{1},d_{2},\cdots,d_{N_{\mathrm{obs}}}\}$ from $N_{\mathrm{obs}}$ BBH merger events, the total number of events is modeled as an inhomogeneous Poisson process, yielding the likelihood [56, 57, 58, 59]:
[…]
where $N_{\exp}(\Lambda)\equiv\xi(\Lambda)T_{\mathrm{obs}}$ is the expected number of detections over the observation timespan $T_{\mathrm{obs}}$, and $\mathcal{L}(d_{i}|\theta)$ represents the likelihood for the $i$th GW event, obtained by reweighing the individual event posterior with the prior on $\theta$. The term $\xi(\Lambda)$ accounts for selection biases for a population with parameters $\Lambda$:
$\xi(\Lambda)=\int P_{\mathrm{det}}(\theta)\,\mathcal{R}_{\mathrm{pop}}(\theta|\Lambda)\,\mathrm{d}\theta,$ (13)
where $P_{\mathrm{det}}(\theta)$ is the detection probability as a function of source parameters $\theta$ [60].
In practice, $\xi(\Lambda)$ is estimated using simulated injections [61] and approximated through a Monte Carlo integral over found injections [62]:
$\xi(\Lambda)\approx\frac{1}{N_{\mathrm{inj}}}\sum_{j=1}^{N_{\text{found}}}\frac{\mathcal{R}_{\mathrm{pop}}(\theta_{j}|\Lambda)}{p_{\mathrm{draw}}(\theta_{j})},$ (14)
where $N_{\text{inj}}$ is the total number of injections, $N_{\text{found}}$ is the count of successfully detected injections, and $p_{\mathrm{draw}}$ is the probability density function used to draw the injections. Using the posterior samples from each event, the hyper-likelihood in Eq. (12) is computed as:
$\mathcal{L}(\textbf{d}|\Lambda)\propto e^{-N_{\mathrm{exp}}}\prod_{i=1}^{N_{\mathrm{obs}}}T_{\mathrm{obs}}\left\langle\frac{\mathcal{R}_{\mathrm{pop}}(\theta|\Lambda)}{\pi(\theta)}\right\rangle,$ (15)
[…]
Parameter Description Prior
$\nu$ friction parameter. ${\mathcal{U}}(-5,6)$
$R_{0}$ PBH merger rate today in $\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$. ${\mathcal{U}}(0,200)$
$H_{0}\,[\mathrm{km}\,\mathrm{s}^{-1}\mathrm{Mpc}^{-1}]$ Hubble constant. ${\mathcal{U}}(20,120)$
$\Omega_{m}$ Present-day matter density of the Universe. ${\mathcal{U}}(0,1)$
$\gamma$ Slope of the power-law regime for the rate evolution before the point $z_{p}$. ${\mathcal{U}}(0,12)$
$\kappa$ Slope of the power-law regime for the rate evolution after the point $z_{p}$. ${\mathcal{U}}(0,6)$
$z_{p}$ Redshift turning point between the powerlaw regimes with $\gamma$ and $\kappa$. ${\mathcal{U}}(0,4)$
$\alpha$ Spectral index for the power-law of the primary mass distribution. ${\mathcal{U}}(2,8)$
$\beta$ Spectral index for the power-law of the mass ratio distribution. ${\mathcal{U}}(-3,7)$
$m_{\min}\,[M_{\odot}]$ Minimum mass of the power-law component of the primary mass distribution. ${\mathcal{U}}(2,7)$
$m_{\max}\,[M_{\odot}]$ Maximum mass of the power-law component of the primary mass distribution. ${\mathcal{U}}(50,200)$
$\lambda_{g}$ Fraction of the model in the Gaussian component. ${\mathcal{U}}(0,0.2)$
$\mu_{g}\,[M_{\odot}]$ Mean of the Gaussian component in the primary mass distribution. ${\mathcal{U}}(20,50)$
$\sigma_{g}\,[M_{\odot}]$ Width of the Gaussian component in the primary mass distribution. ${\mathcal{U}}(0.4,10)$
$\delta_{m}\,[M_{\odot}]$ Range of mass tapering at the lower end of the mass distribution. ${\mathcal{U}}(0,10)$
Table 1: Parameters and their prior distributions used in the Bayesian parameter estimations.
The merger rate $\mathcal{R}_{\mathrm{pop}}$ in Eq. (12) is given by:
$\mathcal{R}_{\mathrm{pop}}=\frac{1}{1+z}\frac{dV_{\mathrm{c}}}{dz}R_{0}\,\pi(z)\,\pi(m_{1})\,\pi(m_{2}|m_{1}),$ (16)
where $R_{0}$ is the local merger rate, $dV_{\mathrm{c}}/dz$ is the differential comoving volume, and $1/(1+z)$ accounts for time dilation between the source and detector frames [63]. The redshift distribution $\pi(z)$ is modeled as a double power-law [63, 64]:
$\pi(z|\gamma,\kappa,z_{\mathrm{p}})=\left[1+\left(1+z_{\mathrm{p}}\right)^{-\gamma-\kappa}\right]\frac{(1+z)^{\gamma}}{1+\left[(1+z)/(1+z_{\mathrm{p}})\right]^{\gamma+\kappa}},$ (17)
with slopes $\gamma$ and $\kappa$, and a characteristic redshift $z_{\mathrm{p}}$. This parameterization is motivated by the potential correlation between the binary formation rate and the star formation rate [65].
The primary mass distribution $\pi(m_{1})$ follows a “Power Law + Peak” model [62], combining a truncated power-law $\mathcal{P}$ with slope $\alpha$ and a Gaussian component $\mathcal{G}$ with mean $\mu_{\mathrm{g}}$ and standard deviation $\sigma_{\mathrm{g}}$. The mixing fraction between the two components is governed by $\lambda_{\mathrm{g}}$. A sigmoid-like window function $S$ ensures a smooth transition at the lower mass cutoff $m_{\min}$ over a scale $\delta_{m}$. The secondary mass distribution $\pi(m_{2})$ is modeled as a truncated power-law with slope $\beta$ between $m_{\min}$ and the primary mass $m_{1}$ [62].
This comprehensive modeling approach allows for a detailed characterization of the merger rate, redshift distribution, and mass distributions of BBH systems, enabling a thorough investigation of their properties and potential origins. The inclusion of the redshift distribution and its potential link to the star formation rate provides insights into the formation channels and evolutionary history of these systems [63, 64]. The mass distributions, with the power-law and Gaussian components, capture the observed features in the black hole mass spectrum and aid in understanding the formation mechanisms and possible existence of a mass gap [62].
Results and Discussion. The prior distributions for the model parameters are summarized in Table 1. The likelihood function is evaluated using the ICAROGW package [59]. The parameter space is explored using the dynesty nested sampling algorithm [66], which is called from the Bilby package [67, 68]. We report the Bayesian posterior distribution of the parameter $\nu$ in Fig. 1. The median value of the parameter, together with its $90\%$ equal-tailed credible interval, is $\nu=0.5^{+3.5}_{-2.6}$. Compared to the constraint from the bright siren of GW170817, which yielded $-75.3\leq\nu\leq 78.4$ [12], the constraint from the spectral siren method using GWTC-3 is improved by an order of magnitude. This improvement can be attributed to the higher redshifts of the BBH mergers in GWTC-3, with the farthest BBH reaching a redshift of $z\sim 1$ [62], while the redshift of GW170817 is only $z\sim 0.008$ [24].
Figure 1: Marginalized posterior distribution for the $\nu$ parameter constrained by the GWTC-3 catalog. The red dashed line indicates the case where $\nu=0$, corresponding to the prediction from GR.
The improved constraint on $\nu$ has important implications for theories of modified gravity and cosmology. A value of $\nu$ consistent with zero, as suggested by our results, indicates that the propagation of GWs is not significantly modified compared to the predictions of GR. This can place constraints on alternative theories of gravity that predict a time-varying Planck mass or a running of the gravitational coupling strength [40, 69, 10]. Such theories, which include scalar-tensor theories like Horndeski and beyond-Horndeski models, are motivated by the desire to explain the accelerated expansion of the Universe without invoking a cosmological constant [6, 7, 70, 71].
Our results also have implications for the physics of the early Universe. A nonzero value of $\nu$ could affect the propagation of primordial GWs generated during inflation [72, 11, 73]. The improved constraint on $\nu$ can be used to refine predictions for the spectrum of the stochastic GW background and its potential detectability by future GW observatories like LISA [74, 75, 76].
The spectral siren method employed in this letter relies on the assumption that the BBH population follows a known redshift distribution, which is inferred from galaxy catalogs and cosmological simulations [63, 77]. Uncertainties in this assumed redshift distribution can propagate to the constraint on $\nu$. In future work, it will be important to quantify the impact of these uncertainties and to explore alternative methods for inferring the redshift distribution of BBHs [78, 79, 80].
Another potential source of systematic uncertainty is the calibration of the GW detectors. The strain calibration uncertainty can affect the measured luminosity distances and, consequently, the inferred constraint on $\nu$ [81, 82]. While the calibration uncertainty is sub-dominant compared to the statistical uncertainty for GWTC-3, it may become more significant as the number of detected events increases with future observing runs [83, 84].
In conclusion, we have used the spectral siren method to constrain the parameter $\nu$, which characterizes the nonstandard propagation of GWs, using the GWTC-3 catalog of BBH mergers. Our results provide an order-of-magnitude improvement over previous constraints from the bright siren GW170817, placing limits on modified gravity theories and the physics of the early Universe. As the sensitivity of GW detectors improves and the number of detected events increases, the spectral siren method will continue to provide a powerful tool for testing GR and probing the nature of gravity on cosmological scales.
Acknowledgments. We are very grateful to Shuxun Tian and Tao Zhu for the fruitful discussions that greatly improve the manuscript. ZCC is supported by the National Natural Science Foundation of China (Grant No. 12247176 and No. 12247112) and the innovative research group of Hunan Province under Grant No. 2024JJ1006. LL is supported by the National Natural Science Foundation of China (Grant No. 12247112 and No. 12247176) and the China Postdoctoral Science Foundation Fellowship No. 2023M730300.